# Impact of Dataset Properties on Membership Inference Vulnerability of Deep Transfer Learning

### v3

@@ C.4 Proof of Theorem 4
$\leq\Pr_{t}\left(|t|\geq F_{|t|}^{-1}(1-\alpha)+\frac{\psi(x,C)}{\sqrt{S}s}\right)-\Pr_{t}(|t|\leq F_{|t|}^{-1}(1-\alpha))$ (A122)
$\approx p_{|t|}(F_{|t|}^{-1}(1-\alpha))\frac{\psi(x,C)}{\sqrt{S}}.$ (A123)
Since $|t|$ follows the standard folded normal distribution with mean 00 and scale $1$, we have
$p_{|t|}(F_{|t|}^{-1}(1-\alpha))=\frac{1}{\sqrt{\pi/2}}\exp\left(-\frac{1}{2}F_{|t|}^{-1}(1-\alpha)^{2}\right).$ (A124)
Hence we obtain
@@
$\log(\textsc{tpr}_{\mathrm{RMIA}}(x)-\textsc{fpr}_{\mathrm{RMIA}}(x))\leq-\frac{1}{2}\log S-\frac{1}{2}F_{|t|}^{-1}(1-\alpha)^{2}+\log\frac{\psi(x,C)}{\sqrt{\pi/2}}.$ (A125)
S FPR=0.2 FPR=0.1 FPR=0.01 FPR=0.001
10 0.000 0.000 0.000 0.000
100 0.000 0.000 0.000 0.000
1000 0.000 0.000 0.041 0.068
10000 0.329 0.353 0.365 0.368
Table A1: The gap in the upper bound for $\textsc{tpr}_{\mathrm{RMIA}}(x)$ in Equation A116 minus the gap in the upper bound for $\textsc{fpr}_{\mathrm{RMIA}}(x)$ in Equation A119. 5 random target examples are sampled from the simplified model distribution with $C=2,s=0.1,d=50$. Simulation is done with 100000 samples of test points $z$, 1000 samples of $t$ and $\gamma=1.0001$. Due to computational constraints, results are not stable especially for small $S$.
C.5 Proof of Corollary 5
See 5
@@ C.5 Proof of Corollary 5
[…]
$=\frac{1}{\sqrt{S}}e^{-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}}\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}\left[\frac{\langle x,x-m_{x}\rangle}{\sqrt{2\pi}||x||s}\right].$ (A129)
Note that here we fixed $\textsc{fpr}_{\mathrm{LiRA}}(x)=\overline{\textsc{fpr}}_{\mathrm{LiRA}}$ for all $x$. Then we obtain
$\log(\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}})\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr})^{2}+\log\left(\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}\left[\frac{\langle x,x-m_{x}\rangle}{\sqrt{2\pi}||x||s}\right]\right).$ (A130)
C.6 Proof of Corollary 6
See 6
Proof.
[…]
$=\frac{1}{\sqrt{S}}e^{-\frac{1}{2}\Phi^{-1}(\alpha)^{2}}\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}\left[\frac{\psi(x,C)}{\sqrt{2\pi}}\right].$ (A134)
Hence we obtain
$\log(\overline{\textsc{tpr}}_{\mathrm{RMIA}}-\overline{\textsc{fpr}}_{\mathrm{RMIA}})\leq-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\alpha)^{2}+\log\left(\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}\left[\frac{\psi(x,C)}{\sqrt{2\pi}}\right]\right).$ (A135)
Appendix D LiRA vulnerability for Small FPR
In Section 3.4 we proved for the simplified model that for LiRA $\log(\textsc{tpr}-\textsc{fpr})\approx-\frac{1}{2}\log S$ ignoring additive constants when fpr is fixed (Corollary 5). In Section 4.3 we observed that the coefficient $\beta_{S}$ for $\log S$ is around $-0.5$ for larger fpr, aligning with the theoretical value. However, for smaller fpr the coefficient is smaller than $-0.5$. To understand this phenomenon, it is important to note that the power-law (Theorem 3 and Corollary 5) only holds for sufficiently large $S$. Thus, it can be explained that the difference of coefficient values $\beta_{S}$ for small and large fpr comes from the small-$S$ regime as follows.
In the proof of Corollary 5 and Theorem 3 the only approximation that could introduce some bias is Equation A84. That is,
$\textsc{tpr}_{\mathrm{LiRA}}(x)-\textsc{fpr}_{\mathrm{LiRA}}(x)$ $=\Pr\left(\eta\leq\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))+r\right)-\Pr\left(\eta\leq\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))\right)$ (A136)
$\approx p_{\eta}\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))\right)r,$ (A137)
where $\eta\sim\mathcal{N}(0,1)$, $r$ is a shift that scales $O(1/\sqrt{S})$ and $p_{\eta}$ is the pdf of $\eta$. Figure A.1 numerically illustrates this approximation. It can be observed that for small $S$ the approximation does not hold, underestimating the true vulnerability. Particularly, this effect is remarkably larger for small $\textsc{fpr}_{\mathrm{LiRA}}(x)$. In other words, the simplified model overestimates the coefficient value $\beta_{S}$ for small $\textsc{fpr}_{\mathrm{LiRA}}(x)$ by underestimating the true vulnerability in the small-$S$ regime.
(a) $\textsc{fpr}_{\mathrm{LiRA}}(x)=0.1$
(b) $\textsc{fpr}_{\mathrm{LiRA}}(x)=0.00001$
Figure A.1: True vs. approximated $\textsc{tpr}_{\mathrm{LiRA}}(x)-\textsc{fpr}_{\mathrm{LiRA}}(x)$ for the simplified model with in-class standard deviation $s=0.1$ and dimension $d=100$. One target example $(x,y_{x})$ is randomly sampled from the underlying distribution. Numbers in the plots are percentages of the approximated vulnerabilities over the true vulnerabilities. Figures illustrate the approximation for different $S$ (a) when $\textsc{fpr}_{\mathrm{LiRA}}(x)=0.1$ and (b) when $\textsc{fpr}_{\mathrm{LiRA}}(x)=0.00001$.
Appendix E Training details
E.1 Parameterization
@@ E.4 Compute resources
Appendix F Additional results
@@
F.1 Additional results for Section 4
@@
F.1.1 Vulnerability as a function of shots
This section displays additional results to Figure 1(a) for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.2, A4 and A5.
@@ F.1.1 Vulnerability as a function of shots
F.1.2 Vulnerability as a function of the number of classes
This section displays additional results to Figure 1(b) for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.3, A6 and A7.
@@ F.1.2 Vulnerability as a function of the number of classes
F.1.3 Data for FiLM and from scratch training
@@
F.1.4 Predicting dataset vulnerability as function of $S$ and $C$
@@
Table A9: Results for fitting Equation A138 with statsmodels Seabold & Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001,0.0001,0.00001\}$. We utilize an ordinary least squares. The test $R^{2}$ assesses the fit to the data of R-50 Head.
coeff. fpr $R^{2}$ test $R^{2}$ coeff. value std. error $t$ $p>|z|$ coeff. $[0.025$ coeff. $0.975]$ $\beta_{S}$ (for $S$) 0.1 0.952 0.907 -0.506 0.011 -44.936 $0$0.529 -0.484 0.01 0.946 0.854 -0.555 0.014 -39.788 $0$0.582 -0.527 0.001 0.930 0.790 -0.627 0.019 -32.722 $0$0.664 -0.589 0.0001 0.852 0.618 -0.741 0.035 -21.467 $0$0.809 -0.673 0.00001 0.837 0.404 -0.836 0.045 -18.690 $0$0.924 -0.748 $\beta_{C}$ (for $C$) 0.1 0.952 0.907 0.090 0.021 4.231 0.000 0.048 0.131 0.01 0.946 0.854 0.182 0.026 6.960 0.000 0.131 0.234 0.001 0.930 0.790 0.300 0.036 8.335 0.000 0.229 0.371 0.0001 0.852 0.618 0.363 0.065 5.616 0.000 0.236 0.491 0.00001 0.837 0.404 0.569 0.085 6.655 0.000 0.400 0.737 $\beta_{0}$ (intercept) 0.1 0.952 0.907 0.314 0.045 6.953 0.000 0.225 0.402 0.01 0.946 0.854 0.083 0.056 1.491 0.137 -0.027 0.193 0.001 0.930 0.790 -0.173 0.077 -2.261 0.025 -0.324 -0.022 0.0001 0.852 0.618 -0.303 0.138 -2.202 0.029 -0.575 -0.032 0.00001 0.837 0.404 -0.615 0.180 -3.414 0.001 -0.970 -0.260
@@
[…]
F.2 Simpler variant of the prediction model
@@
Table A10: Results for fitting Equation A138 with statsmodels Seabold & Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001,0.0001,0.00001\}$. We utilize an ordinary least squares. The test $R^{2}$ assesses the fit to the data of R-50 Head.
coeff. fpr $R^{2}$ test $R^{2}$ coeff. value std. error $t$ $p>|z|$ coeff. $[0.025$ coeff. $0.975]$ $\beta_{S}$ (for $S$) 0.1 0.908 0.764 -0.248 0.008 -30.976 $0$0.264 -0.233 0.01 0.940 0.761 -0.416 0.011 -36.706 $0$0.438 -0.393 0.001 0.931 0.782 -0.553 0.017 -32.507 $0$0.586 -0.519 0.0001 0.865 0.628 -0.697 0.031 -22.274 $0$0.758 -0.635 0.00001 0.862 0.400 -0.802 0.040 -20.311 $0$0.880 -0.725 $\beta_{C}$ (for $C$) 0.1 0.908 0.764 0.060 0.015 3.955 0.000 0.030 0.089 0.01 0.940 0.761 0.169 0.021 7.941 0.000 0.127 0.211 0.001 0.931 0.782 0.297 0.032 9.303 0.000 0.234 0.360 0.0001 0.865 0.628 0.371 0.059 6.328 0.000 0.255 0.486 0.00001 0.862 0.400 0.580 0.076 7.679 0.000 0.431 0.729 $\beta_{0}$ (intercept) 0.1 0.908 0.764 0.029 0.032 0.913 0.362 -0.034 0.093 0.01 0.940 0.761 -0.118 0.045 -2.613 0.010 -0.208 -0.029 0.001 0.931 0.782 -0.295 0.068 -4.345 $0$0.429 -0.161 0.0001 0.865 0.628 -0.387 0.125 -3.104 0.002 -0.633 -0.141 0.00001 0.862 0.400 -0.683 0.159 -4.288 $0$0.996 -0.369
@@
F.3 Empirical results for RMIA
Figures A.6, A.7 and A.8 report additional results for RMIA Zarifzadeh et al. (2024).

### v4

@@ B.3 RMIA power-law
$\kappa_{\textsc{tpr}}(1/\sqrt{S})-\kappa_{\textsc{fpr}}(1/\sqrt{S})\approx p_{\hat{\lambda}_{\bm{z}}}\left(\sqrt{S}\cdot\frac{1}{\sqrt{S}}\right)\frac{c_{0}}{\sqrt{S}}=p_{\hat{\lambda}_{\bm{z}}}(1)\frac{c_{0}}{\sqrt{S}}.$ (A148)
This quantity scales as $O(1/\sqrt{S})$. Therefore,
$\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})-\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})$ $=\Pr_{t}(\kappa_{\textsc{tpr}}(1/\sqrt{S})\geq\tau)-\Pr_{t}(\kappa_{\textsc{fpr}}(1/\sqrt{S})\geq\tau)$ (A149)
$\approx p_{\kappa_{\textsc{fpr}}(1/\sqrt{S})}(\tau)(\kappa_{\textsc{fpr}}(1/\sqrt{S})-\kappa_{\textsc{tpr}}(1/\sqrt{S}))$ (A150)
$\approx p_{\kappa_{\textsc{fpr}}(1/\sqrt{S})}(\tau)p_{\hat{\lambda}_{\bm{z}}}(1)\frac{c_{0}}{\sqrt{S}}.$ (A151)
Note that $\kappa_{\textsc{fpr}}(1/\sqrt{S})$ and, consequently, $\tau$ are independent of $S$ for large enough $S$ since
$\kappa_{\textsc{fpr}}(1/\sqrt{S})=\Pr_{\bm{z}}\left(\lambda_{\bm{z}}\geq\frac{1}{\sqrt{S}}\right)=\Pr_{\bm{z}}\left(\frac{\hat{\lambda}_{\bm{z}}}{\sqrt{S}}\geq\frac{1}{\sqrt{S}}\right)=\Pr_{\bm{z}}(\hat{\lambda}_{\bm{z}}\geq 1).$ (A152)
Hence $\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})-\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})=O(1/\sqrt{S})$. On the other hand, from Equations A139 and A142 we have for sufficiently large $S$
$\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]}{\log\gamma}\geq\tau\right)-\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]}{\log\gamma}+\frac{\mathbb{E}_{\bm{z}}[|q|]}{\log\gamma}\geq\tau\right)$ (A153)
$=$ $\Pr_{Z}\left(\sqrt{S}\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]\geq\tau\right)-\Pr_{Z}\left(\sqrt{S}\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]+\frac{c_{1}}{\sqrt{S}}\geq\tau\right)$ (A154)
$=$ $\Pr_{Z}\left(\mathbb{E}_{\bm{z}}[|\hat{\lambda}_{\bm{z}}|]\geq\tau\right)-\Pr_{Z}\left(\mathbb{E}_{\bm{z}}[|\hat{\lambda}_{\bm{z}}|]+\frac{c_{1}}{\sqrt{S}}\geq\tau\right)$ (A155)
$\approx$ $p_{\mathbb{E}_{\bm{z}}[|\hat{\lambda}_{\bm{z}}|]}(\tau)\frac{c_{1}}{\sqrt{S}},$ (A156)
where $c_{1}$ is some constant. Since for large $S$, $\mathbb{E}_{\bm{z}}[|\hat{\lambda}_{\bm{z}}|]$ and $\tau$ are independent of $S$, Equation A156 scales as $O(1/\sqrt{S})$.
Now let
$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})$ $=\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]+\mathbb{E}_{\bm{z}}[|q|]}{\log\gamma}\geq\tau\right)-v_{\textsc{tpr}}$ (A157)
$\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})$ $=\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]}{\log\gamma}\geq\tau\right)-v_{\textsc{fpr}}$ (A158)
for some $v_{\textsc{tpr}},v_{\textsc{fpr}}\geq 0$ which evaluate the tightness of the bounds. Then we have
$v_{\textsc{tpr}}-v_{\textsc{fpr}}$ (A159)
$=\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]+\mathbb{E}_{\bm{z}}[|q|]}{\log\gamma}\geq\tau\right)-\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]}{\log\gamma}\geq\tau\right)-(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))$ (A160)
$=O(1/\sqrt{S}).$ (A161)
Hence we conclude that for sufficiently large $S$, the bound for $\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})$ is as tight as that for $\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})$.
Therefore, noting that $\mathbb{E}_{\bm{z}}[|q|]/\mathbb{E}_{\bm{z}}[|A|]=O(1/\sqrt{S})$, for large $S$ we obtain
$\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})-\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})$ (A162)
$\approx\Pr_{Z}\left(|Z|\geq F_{|Z|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}\right)-\Pr_{Z}\left(|Z|\geq F_{|Z|}^{-1}(1-\alpha)\right)$ (A163)
$\approx p_{|Z|}\left(F_{|Z|}^{-1}(1-\alpha)\right)\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}$ (A164)
$=p_{|Z|}\left(F_{|Z|}^{-1}(1-\alpha)\right)\frac{\Psi(\bm{x},C)}{\sqrt{S}}.$ (A165)
Since $|Z|$ follows the standard folded normal distribution,
$p_{|Z|}\left(F_{|Z|}^{-1}(1-\alpha)\right)=\frac{2}{\sqrt{2\pi}}\exp\left(-\frac{1}{2}F_{|Z|}^{-1}(1-\alpha)^{2}\right).$ (A166)
It follows that
$\log(\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})-\textsc{fpr}_{\mathrm{RMIA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}F_{|Z|}^{-1}(1-\alpha)^{2}+\log\frac{\Psi(\bm{x},C)}{\sqrt{\pi/2}}.$ (A167)
As for the LiRA power-law, bounding $||\bm{x}-m_{\bm{x}}||$ and $||\bm{z}-m_{\bm{z}}||$ will provide a worst-case upper bound for which the power-law holds.
Finally, the per-example RMIA power-law is also extended to the average-case:
Corollary 7 (Average-case RMIA power-law).
@@
$\log(\overline{\textsc{tpr}}_{\mathrm{RMIA}}-\overline{\textsc{fpr}}_{\mathrm{RMIA}})\approx-\frac{1}{2}\log S-\frac{1}{2}F_{|Z|}^{-1}(1-\alpha)^{2}+\log\left(\mathop{\mathbb{E}}_{(\bm{x},y_{\bm{x}})\sim{\mathbb{D}}}\left[\frac{\Psi(\bm{x},C)}{\sqrt{\pi/2}}\right]\right),$ (A168)
where $\alpha\geq\overline{\textsc{fpr}}_{\mathrm{RMIA}}$ and $F_{|Z|}$ is the cdf of the standard folded normal distribution.
@@
$\overline{\textsc{tpr}}_{\mathrm{RMIA}}-\overline{\textsc{fpr}}_{\mathrm{RMIA}}$ $=\int_{\mathscr{D}}p(\bm{x})(\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})-\textsc{fpr}_{\mathrm{RMIA}}(\bm{x}))\mathrm{d}\bm{x}$ (A169)
$\approx\int_{\mathscr{D}}p(\bm{x})\frac{1}{\sqrt{\pi/2}}e^{-\frac{1}{2}F_{|Z|}^{-1}(1-\alpha)^{2}}\frac{\Psi(\bm{x},C)}{\sqrt{S}}\mathrm{d}\bm{x}$ (A170)
$=\frac{1}{\sqrt{\pi/2}}e^{-\frac{1}{2}F_{|Z|}^{-1}(1-\alpha)^{2}}\int_{\mathscr{D}}p(\bm{x})\frac{\Psi(\bm{x},C)}{\sqrt{S}}\mathrm{d}\bm{x}.$ (A171)
$=\frac{1}{\sqrt{S}}e^{-\frac{1}{2}F_{|Z|}^{-1}(1-\alpha)^{2}}\mathop{\mathbb{E}}_{(\bm{x},y_{\bm{x}})\sim{\mathbb{D}}}\left[\frac{\Psi(\bm{x},C)}{\sqrt{\pi/2}}\right],$ (A172)
where $p(\bm{x})$ is the density of ${\mathbb{D}}$ at $(\bm{x},y_{\bm{x}})$, and $\mathscr{D}$ is the data domain. Hence we obtain
$\log(\overline{\textsc{tpr}}_{\mathrm{RMIA}}-\overline{\textsc{fpr}}_{\mathrm{RMIA}})\approx-\frac{1}{2}\log S-\frac{1}{2}F_{|Z|}^{-1}(1-\alpha)^{2}+\log\left(\mathop{\mathbb{E}}_{(\bm{x},y_{\bm{x}})\sim{\mathbb{D}}}\left[\frac{\Psi(\bm{x},C)}{\sqrt{\pi/2}}\right]\right).$ (A173)
Appendix C Training details
C.1 Parameterization
@@ C.4 Compute resources
Appendix D Additional results
@@
D.1 Additional results for Section 4
@@
D.1.1 Vulnerability as a function of shots
This section displays additional results to Figure 1 for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.1, A3 and A4.
@@ D.1.1 Vulnerability as a function of shots
D.1.2 Vulnerability as a function of the number of classes
This section displays additional results to Figure 2 for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.2, A5 and A6.
@@ D.1.2 Vulnerability as a function of the number of classes
D.1.3 Data for FiLM and from scratch training
@@
D.1.4 Predicting dataset vulnerability as function of $S$ and $C$
@@
Table A8: Results for fitting Equation 13 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001\}$. We utilize an ordinary least squares. The test $R^{2}$ assesses the fit to the data of R-50 Head.
coeff. fpr $R^{2}$ test $R^{2}$ coeff. value std. error $t$ $p>|z|$ coeff. $[0.025$ coeff. $0.975]$ $\beta_{S}$ (for $S$) 0.1 0.952 0.907 -0.506 0.011 -44.936 $0$0.529 -0.484 0.01 0.946 0.854 -0.555 0.014 -39.788 $0$0.582 -0.527 0.001 0.930 0.790 -0.627 0.019 -32.722 $0$0.664 -0.589 $\beta_{C}$ (for $C$) 0.1 0.952 0.907 0.090 0.021 4.231 0.000 0.048 0.131 0.01 0.946 0.854 0.182 0.026 6.960 0.000 0.131 0.234 0.001 0.930 0.790 0.300 0.036 8.335 0.000 0.229 0.371 $\beta_{0}$ (intercept) 0.1 0.952 0.907 0.314 0.045 6.953 0.000 0.225 0.402 0.01 0.946 0.854 0.083 0.056 1.491 0.137 -0.027 0.193 0.001 0.930 0.790 -0.173 0.077 -2.261 0.025 -0.324 -0.022
@@
D.2 Simpler variant of the prediction model
@@
Table A9: Results for fitting Equation A174 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001\}$. We utilize an ordinary least squares. The test $R^{2}$ assesses the fit to the data of R-50 Head.
coeff. fpr $R^{2}$ test $R^{2}$ coeff. value std. error $t$ $p>|z|$ coeff. $[0.025$ coeff. $0.975]$ $\beta_{S}$ (for $S$) 0.1 0.908 0.764 -0.248 0.008 -30.976 $0$0.264 -0.233 0.01 0.940 0.761 -0.416 0.011 -36.706 $0$0.438 -0.393 0.001 0.931 0.782 -0.553 0.017 -32.507 $0$0.586 -0.519 $\beta_{C}$ (for $C$) 0.1 0.908 0.764 0.060 0.015 3.955 0.000 0.030 0.089 0.01 0.940 0.761 0.169 0.021 7.941 0.000 0.127 0.211 0.001 0.931 0.782 0.297 0.032 9.303 0.000 0.234 0.360 $\beta_{0}$ (intercept) 0.1 0.908 0.764 0.029 0.032 0.913 0.362 -0.034 0.093 0.01 0.940 0.761 -0.118 0.045 -2.613 0.010 -0.208 -0.029 0.001 0.931 0.782 -0.295 0.068 -4.345 $0$0.429 -0.161
@@
D.3 Empirical results for RMIA
Figures A.5, A.6 and A.7 report additional results for RMIA Zarifzadeh et al. (2024).